OpenStudy (anonymous):

What is the rate of change for the equation 5x+3y=-2

OpenStudy (anonymous):

Solve the equation for y, and then look at the slope.

OpenStudy (anonymous):

rate of change = dy/dx; first step is to solve for y 3y=-2-5x y=-2/3 -5/3x dy/dx = -5/3

OpenStudy (anonymous):

why and how is it -5/3